1. A stone is thrown straight up at a speed of 12 m / s.

a. After how long does the stone turn down again?

b. How high does the stone reach before it turns?


1
Expert's answer
2021-04-15T10:45:09-0400

We have given the initial velocity "u = 12ms^{-1}"

a.) Let the time taken by the stone turn down again = t

Displacement s =0,

We know,

"s= ut+\dfrac{1}{2}gt^2"


"0 = (-12)t+ \dfrac{1}{2}(10)t^2"


"t = \dfrac{12\times2}{10}"


"t = 2.4s"


b.)Height up to the stone reach before it turns = maximum height it achieves

Max. height can be calculated using third equation of motion,


"v^2-u^2 = 2gh"


"(0)^2-(12)^2 = 2\times -10 \times h"


"h= 7.2m"


"h= H_{max} = 7.2m"